OpenStudy (anonymous):

If f (x) = (4 sin x + 2)^(1/2), f' (0) = ???

OpenStudy (anonymous):

Use u^n du rule here. u = 4sinx + 2 f(x) = (4sinx + 2)^1/2 Chain rule f'(x) = (1/2)(4cosx)(4sinx+2)^(-1/2) f'(x)=2cosx(4sinx+2)^-1/2

OpenStudy (isaiah.feynman):

To find the derivative of a square root use this fast trick! All you have to do is just find the derivative of the expression under the root!